OpenStudy (anonymous):

you borrow $1000 from a friend and promise to pay back $1390 in 2 years. what simple interest rate, to the nearest tenth of a percent will you pay

OpenStudy (anonymous):

@DanJS

OpenStudy (texaschic101):

I = PRT <-- our formula (Interest = Principle* Rate*Time) The interest paid back was (1390 - 1000 = 390) I = 390 P = 1000 T = 2 Now we sub 390 = 1000(R)(2) 390 = 2000R 390/2000 = R 0.195 = R 19.5% = R Interest rate 19.5%

OpenStudy (anonymous):

THANK YOU

OpenStudy (texaschic101):

glad to help :)